# Bridging the Rossby number gap in rapidly rotating thermal convection

### v1

@@ Discussion
As described in detail in Ref.  (?), the power behind the Rescaled incompressible Navier-Stokes Equations (RiNSE) lies in their use of rotation-influenced characteristic scales for nondimensionalisation. For convenience, we introduce the small parameter $\varepsilon\equiv{\rm Ek}^{1/3}=\left(\nu/2\Omega H^{2}\right)^{1/3}$ and measure lengths in units of the layer depth $H$ in the vertical, but in units of $\ell_{*}=\varepsilon H$ in the horizontal, so that $(X,Y,Z)=(\varepsilon H\widetilde{x},\varepsilon H\widetilde{y},H\widetilde{z})$. In these coordinates, the Laplace operator becomes $\Delta=\left(\varepsilon^{-2}H^{-2}\right)\widetilde{\nabla}_{\varepsilon}^{2}$ where we have introduced the anisotropic diffusion operator: ~$\widetilde{\nabla}_{\varepsilon}^{2}=\partial_{\tilde{x}\tilde{x}}+\partial_{\tilde{y}\tilde{y}}+\varepsilon^{2}\partial_{\tilde{z}\tilde{z}}$. All tildes will be dropped in the following. Time is non-dimensionalized using a diffusive time-scale based on the horizontal scale rather than the layer depth alone: $t_{*}=\ell_{*}^{2}/\nu=\varepsilon^{2}H^{2}/\nu$. Accordingly, the velocity scale is $u_{*}=\ell_{*}/t_{*}=\nu/(\varepsilon H)$, giving the nondimensional velocity $\boldsymbol{u}=\boldsymbol{U}/u_{*}\equiv(u,v,w)$. This isotropic velocity scale, combined with our anisotropic coordinate system, results in horizontally-dominated material derivatives:
@@
Reduced non-hydrostatic quasi-geostrophic equations
In the limit of rapid rotation ${\rm Ek},{\rm Ro}\to 0$, one can derive a reduced set of governing equations for the flow (presented below) in a rigorous and controlled way using asymptotic methods  (?, ?, ?, ?). Here we emphasize, in a quick sketch, how the rescaled equations (8) naturally lead to the reduced equations when the formal limit $\varepsilon\to 0$ is considered. Upon inspection, if the scaling employed for deriving the rescaled equations holds and all dimensionless variables remain $O(1)$, the leading order terms at order $O(\varepsilon^{-1})$ in equation (8a) indicate geostrophic balance:
@@
[…]
@@
Equations (13) form the reduced Non-Hydrostatic Quasi-Geostrophic Equations (NHQGE), which were previously derived using a multiple-scale asymptotic expansion  (?).
@@
The Ekman number dependence that remains in the governing equation (13d) for the mean temperature reflects the fact that the timescale $\tau=\varepsilon^{2}t$ associated with the evolution of $\Theta$ is much slower than the convective timescale. Previous studies  (?, ?, ?) have established that, in the limit of horizontally extended domains where a large number of convective cells or plumes contribute to the horizontal spatial averaging, this term can be omitted for the computation of global diagnostics of a statistically stationary state $\partial_{t}\overline{\Theta}\approx 0$. This assumption can be justified based on the following rationale: as the horizontal extent $L_{\perp}=L_{x}=L_{y}$ of the domain increases, so does the number of independent plumes $N_{p}$ that feed back into the temperature profile. The central limit theorem then suggests that, in the ergodic limit (time and ensemble averages behave similarly), the amplitude of temporal fluctuations of the vertical temperature profile $\Theta(z)$ should decrease like $1/\sqrt{N_{p}}\approx 1/L_{\perp}$, justifying the neglect of the term $\partial_{t}\Theta$. In the context of the RiNSE formulation, this assumption has been successfully validated and has proved to shorten the transient that leads to the stationary state  (?).
Numerical simulations
@@
We compute solutions to both the rescaled and the reduced equations using the pseudo-spectral code Coral  (?), a flexible platform for solving partial differential equations in a Cartesian domain. We suppose periodicity along the horizontal directions $(x,y)$ and thus decompose the problem variables using a Fourier basis, whereas a Chebyshev expansion is used in the vertical in order to impose boundary conditions on the bounding surfaces $z=0,1$ using a Galerkin basis recombination. Differential operators are discretized using the quasi-inverse method  (?, ?) to optimize memory usage and accuracy. The resolution was varied from $192$ grid points (corresponding to $128$ alias-free modes) at the lowest values of $\rm\widetilde{Ra}$ in each direction, and up to $768$ grid points ($512$ Chebychev modes) in the vertical direction as well as $588$ grid points ($384$ Fourier modes) in the horizontal directions. The time-marching relies on a family of stable semi-implicit Runge-Kutta schemes  (?), specifically the third-order ARS443 scheme.
@@
For some runs presented here, we have used low amplitude noise as the initial condition, and waited for the convective turbulent flow to settle into a statistically stationary state after a transient period. However, for strongly supercritical flow, this transient may prove much more turbulent than the stationary state itself. Hence, in some cases, we resort to an alternative: instead of low amplitude noise, the initial condition for a given run would be chosen as the solution computed at a neighboring position in parameter space. This procedure facilitates the computation of the transient.

### v2

@@ 2.1 Rescaled Navier-Stokes equation
As described in detail by Julien et al. (2024b), the power behind the Rescaled incompressible Navier-Stokes Equations (RiNSE) lies in their use of rotation-influenced characteristic scales for nondimensionalisation. For convenience, we introduce the small parameter $\varepsilon\equiv{\rm Ek}^{1/3}=\left(\nu/2\Omega H^{2}\right)^{1/3}$ and measure lengths in units of the layer depth $H$ in the vertical, but in units of $\ell_{*}=\varepsilon H$ in the horizontal, so that $(X,Y,Z)=(\varepsilon H\widetilde{x},\varepsilon H\widetilde{y},H\widetilde{z})$. In these coordinates, the Laplace operator becomes $\Delta=\left(\varepsilon^{-2}H^{-2}\right)\widetilde{\nabla}_{\varepsilon}^{2}$ where we have introduced the anisotropic diffusion operator: ~$\widetilde{\nabla}_{\varepsilon}^{2}=\partial_{\tilde{x}\tilde{x}}+\partial_{\tilde{y}\tilde{y}}+\varepsilon^{2}\partial_{\tilde{z}\tilde{z}}$. All tildes will be dropped in the following. Time is nondimensionalized using a diffusive time-scale based on the horizontal scale rather than the layer depth alone: $t_{*}=\ell_{*}^{2}/\nu=\varepsilon^{2}H^{2}/\nu$. Accordingly, the velocity scale is $u_{*}=\ell_{*}/t_{*}=\nu/(\varepsilon H)$, giving the nondimensional velocity $\boldsymbol{u}=\boldsymbol{U}/u_{*}\equiv(u,v,w)$. This isotropic velocity scale, combined with our anisotropic coordinate system, results in horizontally-dominated material derivatives:
@@
2.2 Asymptotically reduced non-hydrostatic quasi-geostrophic equations
In the limit of rapid rotation ${\rm Ek},{\rm Ro}\to 0$, one can derive a reduced set of governing equations for the flow (presented below) in a systematic and controlled way using asymptotic methods (Julien et al., 1998; Sprague et al., 2006; Julien & Knobloch, 2007; Julien et al., 2012b). Here we emphasize, in a quick sketch, how the rescaled equations (6) naturally lead to the reduced equations when the formal limit $\varepsilon\to 0$ is considered. Upon inspection, if the scaling employed for deriving the rescaled equations holds and all dimensionless variables remain $O(1)$, the leading order terms at order $O(\varepsilon^{-1})$ in equation (6a) indicate geostrophic balance:
@@ 2.2 Asymptotically reduced non-hydrostatic quasi-geostrophic equations
where $J[p,w]\equiv\partial_{x}p\partial_{y}w-\partial_{y}p\partial_{x}w$ indicates horizontal advection of $w$ and $\nabla_{\perp}^{2}\equiv\partial_{xx}+\partial_{yy}$ emphasizes that horizontal diffusion dominates. We complement the vertical velocity equation with a governing equation for the streamfunction $p$ obtained by projecting the curl of the momentum equation on the vertical. Using incompressibility, one proves $\mathbf{\widehat{e}}_{z}\boldsymbol{\cdot\nabla}\times\left(\mathbf{\widehat{e}}_{z}\times\boldsymbol{u}\right)=-\varepsilon\partial_{z}w$, so that the leading order contribution of $\mathbf{\widehat{e}}_{z}\boldsymbol{\cdot\nabla}\times$ (6a) yields:
@@
Equations (11) form the reduced Non-Hydrostatic Quasi-Geostrophic (NHQG) equations, previously derived using a multiple-scale asymptotic expansion (Julien et al., 1998; Sprague et al., 2006).
@@
The Ekman number dependence that remains in the governing equation (11d) for the mean temperature reflects the fact that the timescale $\tau=\varepsilon^{2}t$ associated with the evolution of $\Theta$ is much slower than the convective timescale. Previous studies (Sprague et al., 2006; Julien et al., 2012b, 1998) have established that, in the limit of horizontally extended domains where a large number of convective cells or plumes contribute to the horizontal spatial averaging, this term can be omitted for the computation of global diagnostics of a statistically stationary state $\partial_{t}\overline{\Theta}\approx 0$. This assumption can be justified based on the following rationale: as the horizontal extent $L_{\perp}=L_{x}=L_{y}$ of the domain increases, so does the number of independent plumes $N_{p}$ that feed back into the temperature profile. The central limit theorem then suggests that, in the ergodic limit (time and ensemble averages behave similarly), the amplitude of temporal fluctuations of the vertical temperature profile $\Theta(z)$ should decrease like $1/\sqrt{N_{p}}\approx 1/L_{\perp}$, justifying the neglect of the term $\partial_{t}\Theta$. In the context of the RiNSE formulation, this assumption has been successfully validated and has proved to shorten the transient that leads to the stationary state (Julien et al., 2024b).
2.3 Details on the numerical method
@@
We compute solutions to both the rescaled and the reduced equations using the pseudo-spectral code Coral (Miquel, 2021), a flexible platform for solving partial differential equations in a Cartesian domain. We suppose periodicity along the horizontal directions $(x,y)$ and thus decompose the problem variables using a Fourier basis, whereas a Chebyshev expansion is used in the vertical in order to impose boundary conditions on the bounding surfaces $z=0,1$ using a Galerkin basis recombination. Differential operators are discretized using the quasi-inverse method (Clenshaw, 1957; Greengard, 1991; Julien & Watson, 2009) to optimize memory usage and accuracy. The resolution was varied from $192$ grid points in each direction (corresponding to $128$ alias-free modes) at the lowest values of $\rm\widetilde{Ra}$, up to $768$ grid points ($512$ Chebychev modes) in the vertical direction and $576$ grid points ($384$ Fourier modes) in the horizontal directions at larger $\rm\widetilde{Ra}$. The time-marching relies on the family of stable semi-implicit Runge-Kutta schemes introduced by Ascher et al. (1997), specifically their third-order ARS443 scheme.
@@
For some of the runs presented here, we have used low amplitude noise as the initial condition, and waited for the convective turbulent flow to settle into a statistically stationary state after a transient period. However, for strongly supercritical flow, this transient may prove much more turbulent than the stationary state itself. Hence, in some cases, we resort to an alternative: instead of low amplitude noise, the initial condition for a given run was taken from the solution computed at a neighboring position in parameter space. This procedure facilitates the computation of the transient.